OpenStudy (anonymous):

What is the equation of the line y= -3/4x +2/3 in standard form? Use integer coefficients.

OpenStudy (anonymous):

standard form : Ax + By = C non standard form = By = Ax +C try to work it out

OpenStudy (anonymous):

\[y = - \frac{ 3 }{ 4 } x + \frac{ 2 }{ 3 }\] \[y = \frac{ -9x +8 }{ 12 }\] \[9x + 12y = 8\]

OpenStudy (anonymous):

thank you! (: